OpenStudy (anonymous):

-4÷(-1/3)×3×(-1/6)-(-9)

OpenStudy (anonymous):

Work through it step by step, use PEMDAS and watch out for double negatives.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

-1/3x3=-1 -1x(-1/6)=1/6 1/6-(-9)=55/6 -4/55/6=-24/55